# Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs

### v4

@@
Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs
Maxim Zhelnin zhelninmax@gmail.com MWS AIMoscowRussia , Dmitry Redko dmitryredko444@gmail.com Applied AIMoscowRussia , Daniil Volkov danvol121@gmail.com Applied AIMoscowRussia , Anna Volodkevich volodkanna@yandex.ru Sber AI Lab, Applied AIMoscowRussia , Petr Sokerin sokerinpo@gmail.com Applied AIMoscowRussia , Valeriy Shevchenko escape756@gmail.com IVIMoscowRussia , Egor Shvetsov e.shvetsov@applied-ai.ru Applied AIMoscowRussia , Alexey Vasilev alexxl.vasilev@yandex.ru Sber AI LabMoscowRussia , Darya Denisova duny.explorer@gmail.com Sber AI LabMoscowRussia , Ruslan Izmailov lord.rik@yandex.ru SberMoscowRussia and Alexey Zaytsev likzet@gmail.com Applied AIMoscowRussia
Sequential Recommendation, Cross-Entropy Loss Optimization, GPU Memory Constraints, Negative Sampling, Training Efficiency, Softmax Saturation, Memory-Efficient Training
$c$s: Information systems Recommender systems
@@ 1. Introduction
User actions can be modeled sequentially, conditioning on past actions in diverse scenarios, ranging from a mechanical series of actions in a banking app (boka2024survey) to a mood- and neighborhood-influenced selection of music tracks (krause2014music). Consequently, Sequential Recommender Systems (SRS), which model user interactions as sequences, have become essential components of modern recommendation pipelines. SRS effectively incorporates ideas from natural language processing (NLP) through attention-based architectures such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec), as well as the cross-entropy loss function (klenitskiy2024does; sun2019bert4rec; mezentsev2024scalable).
@@
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where $bs$ is the batch size, $sl$ sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (wijmans2024cut). However, in SRS, $|V|$ can reach millions of items (zivic2024scaling), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (touvron2023llama), QWEN (bai2023qwen), and Gemma-7B (team2403gemma), respectively.
@@
Dataset BCE CE CE- CCE CCE-
Megamarket 0.0017 0.0052 0.0117 0.0181 0.0168
Zvuk 0.0046 0.0797 0.0675 0.0940 0.0852
30Music 0.0004 0.0891 0.0939 0.1393 0.1436
Gowalla 0.0197 0.0225 0.0278 0.0241 0.0279
Beauty 0.0094 0.0114 0.0123 0.0104 0.0116
ML-20M 0.0557 0.0510 0.0513 0.0511 0.0523
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (kang2018self) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(sun2019bert4rec; klenitskiy2023turning). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (klenitskiy2023turning; mezentsev2024scalable). Instead of scoring all items during training, a subset of negative samples ($ns$) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing $ns$ in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (di2024theoretical; klenitskiy2023turning; guo2024scaling). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (mezentsev2024scalable; klenitskiy2023turning) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(prakash2024evaluating). This creates an open question of whether practitioners should maximize $ns$ given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
Training Speed
Although sampling strategies such as $\mathrm{CE}^{-}$ reduce memory consumption, they also improve training speed by requiring fewer operations, partially addressing the problem of frequent retraining SRS. However, there are other approaches for faster training: quantization (wortsman2023stable), activation sparsification (hu2024accelerating), and finally gradient sparsification (alistarh2018convergence). The latter one is usually overlooked, even though gradients usually require twice as many computations as the forward path. In our work, we demonstrate that gradient sparsification is especially suitable for SRS, provide a formal analysis in Section 3.2, and study how sparse gradients accelerate model training and their impact on accuracy degradation. We will refer to it as our second research question Q2.
@@
Computational efficiency critically depends on the co-design of architecture, hardware, and low-level implementations. While hardware is often fixed, efficiency can be optimized through architectures better aligned with underlying hardware (shvetsov2024quantnas) or specialized low-level optimizations. Though highly specific, such optimizations yield significant computational benefits. Examples include FlashAttention (dao2022flashattention) for efficient attention computation and Liger kernels (hsu2024liger) for optimized CE computation. Among them, Cut-Cross Entropy (CCE) was recently developed for LLMs; it helps to alleviate memory bottlenecks during training when CEloss is employed. However, until now, the CCE has remained unexplored for SRS models. Furthermore, our analysis reveals the importance of negative sampling, which is not supported by the original CCE. Development and analysis of CCE with negative sampling is our third research question - Q3, more precisely: Can we design a CCE alternative that is faster and more memory efficient for SRS?
@@ 2. Related work
Cross-Entropy and sampled CE for SRS. Sequential recommendation has seen significant advancements in recent years. One pivotal observation from  (klenitskiy2023turning) demonstrated that SASRec’s performance improves when trained with cross-entropy loss instead of binary cross-entropy. Unfortunately, computing cross-entropy over all items is computationally prohibitive. To address this, researchers adopted sampled cross-entropy loss, which uses a subset of negative items. This approach was theoretically justified in  (di2024theoretical), showing that increasing negative samples enhances performance. However, negative sampling has non-linear performance scaling. Despite its benefits, negative sampling amplifies popularity bias, as popular items are more likely to be sampled  (prakash2024evaluating). Consequently, model performance becomes highly dependent on dataset distribution and sampling strategy. Furthermore,  (prakash2024evaluating) highlights that performance does not scale linearly with the number of negative samples, which complicates the selection of hyperparameters. To address this problem, we perform a careful and extensive study of memory-dependent hyperparameter analysis for optimal accuracy-memory tradeoff (Section 5.1). Prior work on scaling for SRS models  (guo2024scaling; zhang2024scaling; zivic2024scaling; xu2025efficient) emphasizes the benefits of scaling sequence length and model depth. However, our experiments reveal that expanding a single factor [^3] Factors include batch size, sequence length, and number of negative samples. yields limited gains, underscoring the need for multi-dimensional scaling. Unlike  (mezentsev2024scalable), our work focuses only on feature-scaling [^4] Feature scaling refers to adjusting batch size, sequence length, etc. effects while keeping model architecture fixed.
@@
Computationally efficient approaches. Advanced sampling techniques for balancing accuracy and memory were proposed in  (mezentsev2024scalable). Liger-Kernel  (hsu2024liger), implemented in the Triton language  (tillet2019triton), addresses memory consumption issues through the use of FusedLinearCrossEntropy (FLCE). While theoretical convergence guarantees for sparse gradient methods were established in  (alistarh2018convergence), it’s application for efficient training remains limited  (chekalina2024sparsegrad; wijmans2024cut). Building on FLCE memory-efficient cross-entropy with saturated gradient filtering beyond numerical precision was proposed (CCE) for LLMs  (wijmans2024cut). Since saturated gradients do not fit numerical precision and turn to zero, their removal does not affect model convergence. In this work, we analyze whether the removal of low-magnitude non-zero gradients would help to accelerate model training with minimal performance degradation.
@@
Possible further directions and trends. Even when isolating experiments to feature scaling, there remains significant flexibility in training strategies. For example, work  (li2022stability) demonstrated that training GPT-like models with short sequences initially and gradually increasing them enables stable, efficient training with larger batch sizes and learning rates. Similarly, authors of  (di2024theoretical) found that starting with fewer negative samples and increasing their count later boosts SRS model performance. Moreover, based on our results, an adaptive threshold to prune gradients can be employed. While these adaptive strategies are promising, we focus on studying simpler non-adaptive approaches to establish a stronger foundation for future research.
@@ 3.1. Cut-Cross-Entropy for RecSys
Our proposed method, $\mathrm{CCE}^{-}$, builds upon CCE, which itself extends the Liger-Kernel (LG) (hsu2024liger). Below, we provide a brief review of both $\mathrm{CCE}^{-}$ and LG, highlighting their key differences. We then focus on how $\mathrm{CCE}^{-}$ introduces specific improvements tailored for sequential recommendation systems.
During training, we deal with two types of memory: GPU memory (High Bandwidth Memory, HBM) and on-chip shared memory (Static Random-Access Memory, SRAM). HBM is large but relatively slow and is used to store data for computations. GPU kernel loads input data from HBM into SRAM to perform computations. SRAM is significantly faster than HBM but has much less capacity. Therefore, only small chunks of input data can be loaded into SRAM at a time. Once computations are complete, the output is saved back to HBM. While there are more details, see (dao2022flashattention),  (wijmans2024cut), this structure shapes the main goal: reduce the amount of data loaded into GPU memory. It can be achieved by minimizing the overall amount of memory used by a model or optimizing transfers between HBM and SRAM.
@@ 3.2. Sparse Gradients in Training with CE loss
In the recommender systems, the objective is to predict the next item from a predefined set $V$ of |V||V|{} items. Similar to the next-token prediction, there is exactly one correct class and |V|1|V|{}-1 incorrect classes. Thus, employing loss functions such as the CE loss provides a suitable estimate of model quality  (di2024theoretical).
@@ 3.3. Cut-Cross-Entropy with Negative Sampling
As demonstrated in (klenitskiy2023turning), negative sampling offers dual advantages: it allows better accuracy on certain datasets and accelerates training compared to full vocabulary utilization. Motivated by this, we extend $\mathrm{CCE}^{-}$ by integrating negative sampling into it.
@@
To achieve a memory-efficient computation for this term, we developed a specialized kernel that fuses these operations. For this purpose, we adopt an approach for the online softmax computation. As a result, all intermediate operations are carried out on-chip using shared memory, while only the log-sum-exp vector is stored in the global memory (milakov2018onlinesoftmax; hsuliger).
@@
During the forward pass, the output embedding $\mathbf{E}$ is divided into the blocks $E_{n,D}$ of size $N_{B}\times D$. Since the inner dimension $D$ in transformer models for RecSys is typically not large (in  (kang2018self), $D<50$; in  (klenitskiy2023turning)(klenitskiy2024does), $D=64$; in  (guo2024scaling), $D<400$ ), division is not performed along the $D$ dimension. The matrix of the linear classifier $\mathbf{C}$ is loaded inside a for-loop by columns, determined by the block of indices $\mathbf{Inds}_{n,j}$ of size $N_{B}$.
@@ 4.1. Datasets
We use six datasets: 30Music  (turrin201530music), Zvuk  (shevchenko2024variability), MovieLens-20M (ML 20M)  (harper2015movielens), MegaMarket (shevchenko2024variability), Beauty  (mcauley2015image), and Gowalla  (cheng2013you). Our main focus is on datasets with a large item catalog: Zvuk, MegaMarket, Gowalla, and 30Music, which include more than 300K items. Moreover, Zvuk and MegaMarket are some of the largest open datasets in the recommendation domain  (shevchenko2024variability). MovieLens and Beauty are added because of their popularity, used in more than 14 and 45 papers according to (klenitskiy2024does). For Beauty, we employ its third version, which has the largest item catalog. The characteristics of all datasets after preprocessing are in Table 2.
@@ 4.2. Evaluation Methodology
Global Temporal Cutoff: To ensure alignment with real-world sequential prediction scenarios, we use a global temporal split with the last interaction as a target and ”by user” validation scheme (gusak2025time). We split interactions at the 90% timestamp quantile. All interactions before this cutoff form the training/validation pool, while interactions after constitute the test set. Validation Set: From the training pool, we randomly select 5% of users. For each selected user, their last interaction is withheld for validation. Test Set: For users in the post-cutoff test set, only their last interaction is retained for evaluation, while all previous ones are passed as model input.
@@
Following prior work  (anelli2019discriminative; zhao2022revisiting; shevchenko2024variability), we prioritize NDCG@10 due to its strong correlation with overall ranking quality. Along with it, record other metrics, including Coverage and Surprisal (vasilev2024replay). All metrics are available in our publicly released dataset.
@@
We primarily focus on the SASRec model, as it is more commonly used in practice and frequently outperforms Bert4Rec (klenitskiy2023turning; mezentsev2024scalable). However, to demonstrate the efficiency of CCE and $\mathrm{CCE}^{-}$ for other architectures, we also present the results of applying these loss functions to the training of BERT4Rec in Section 5.6

### v5

@@
Abstract.
@@ 1. Introduction
User actions can be modeled sequentially, conditioning on past actions in diverse scenarios, ranging from a mechanical series of actions in a banking app (Boka et al., 2024) to a mood- and neighborhood-influenced selection of music tracks (Krause et al., 2014). Consequently, Sequential Recommender Systems (SRS), which model user interactions as sequences, have become essential components of modern recommendation pipelines. SRS effectively incorporates ideas from natural language processing (NLP) through attention-based architectures such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), as well as the cross-entropy loss function (Klenitskiy et al., 2024; Sun et al., 2019; Mezentsev et al., 2024).
@@
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where $bs$ is the batch size, $sl$ sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (Wijmans et al., 2024). However, in SRS, $|V|$ can reach millions of items (Zivic et al., 2024), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (Touvron et al., 2023), QWEN (Bai et al., 2023), and Gemma-7B (Team et al., 2024), respectively.
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (Kang and McAuley, 2018) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(Sun et al., 2019; Klenitskiy and Vasilev, 2023). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (Klenitskiy and Vasilev, 2023; Mezentsev et al., 2024). Instead of scoring all items during training, a subset of negative samples ($ns$) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing $ns$ in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (Di Teodoro et al., 2024; Klenitskiy and Vasilev, 2023; Guo et al., 2024). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (Mezentsev et al., 2024; Klenitskiy and Vasilev, 2023) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(Prakash et al., 2024). This creates an open question of whether practitioners should maximize $ns$ given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
Training Speed
Although sampling strategies such as $\mathrm{CE}^{-}$ reduce memory consumption, they also improve training speed by requiring fewer operations, partially addressing the problem of frequent retraining SRS. However, there are other approaches for faster training: quantization (Wortsman et al., 2023), activation sparsification (Hu et al., 2024), and finally gradient sparsification (Alistarh et al., 2018). The latter one is usually overlooked, even though gradients usually require twice as many computations as the forward path. In our work, we demonstrate that gradient sparsification is especially suitable for SRS, provide a formal analysis in Section 3.2, and study how sparse gradients accelerate model training and their impact on accuracy degradation. We will refer to it as our second research question Q2.
@@
Computational efficiency critically depends on the co-design of architecture, hardware, and low-level implementations. While hardware is often fixed, efficiency can be optimized through architectures better aligned with underlying hardware (Shvetsov et al., 2024) or specialized low-level optimizations. Though highly specific, such optimizations yield significant computational benefits. Examples include FlashAttention (Dao et al., 2022) for efficient attention computation and Liger kernels (Hsu et al., 2024a) for optimized CE computation. Among them, Cut-Cross Entropy (CCE) was recently developed for LLMs; it helps to alleviate memory bottlenecks during training when CEloss is employed. However, until now, the CCE has remained unexplored for SRS models. Furthermore, our analysis reveals the importance of negative sampling, which is not supported by the original CCE. Development and analysis of CCE with negative sampling is our third research question - Q3, more precisely: Can we design a CCE alternative that is faster and more memory efficient for SRS?
@@
Dataset BCE CE CE- CCE CCE-
Megamarket 0.0017 0.0052 0.0117 0.0181 0.0168
Zvuk 0.0046 0.0797 0.0675 0.0940 0.0852
30Music 0.0004 0.0891 0.0939 0.1393 0.1436
Gowalla 0.0197 0.0225 0.0278 0.0241 0.0279
Beauty 0.0094 0.0114 0.0123 0.0104 0.0116
ML-20M 0.0557 0.0510 0.0513 0.0511 0.0523
@@ 2. Related work
Cross-Entropy and sampled CE for SRS. Sequential recommendation has seen significant advancements in recent years. One pivotal observation from  (Klenitskiy and Vasilev, 2023) demonstrated that SASRec’s performance improves when trained with cross-entropy loss instead of binary cross-entropy. Unfortunately, computing cross-entropy over all items is computationally prohibitive. To address this, researchers adopted sampled cross-entropy loss, which uses a subset of negative items. This approach was theoretically justified in  (Di Teodoro et al., 2024), showing that increasing negative samples enhances performance. However, negative sampling has non-linear performance scaling. Despite its benefits, negative sampling amplifies popularity bias, as popular items are more likely to be sampled  (Prakash et al., 2024). Consequently, model performance becomes highly dependent on dataset distribution and sampling strategy. Furthermore,  (Prakash et al., 2024) highlights that performance does not scale linearly with the number of negative samples, which complicates the selection of hyperparameters. To address this problem, we perform a careful and extensive study of memory-dependent hyperparameter analysis for optimal accuracy-memory tradeoff (Section 5.1). Prior work on scaling for SRS models  (Guo et al., 2024; Zhang et al., 2024; Zivic et al., 2024; Xu et al., 2025) emphasizes the benefits of scaling sequence length and model depth. However, our experiments reveal that expanding a single factor [^3] Factors include batch size, sequence length, and number of negative samples. yields limited gains, underscoring the need for multi-dimensional scaling. Unlike  (Mezentsev et al., 2024), our work focuses only on feature-scaling [^4] Feature scaling refers to adjusting batch size, sequence length, etc. effects while keeping model architecture fixed.
@@
Computationally efficient approaches. Advanced sampling techniques for balancing accuracy and memory were proposed in  (Mezentsev et al., 2024). Liger-Kernel  (Hsu et al., 2024a), implemented in the Triton language  (Tillet et al., 2019), addresses memory consumption issues through the use of FusedLinearCrossEntropy (FLCE). While theoretical convergence guarantees for sparse gradient methods were established in  (Alistarh et al., 2018), it’s application for efficient training remains limited  (Chekalina et al., 2024; Wijmans et al., 2024). Building on FLCE memory-efficient cross-entropy with saturated gradient filtering beyond numerical precision was proposed (CCE) for LLMs  (Wijmans et al., 2024). Since saturated gradients do not fit numerical precision and turn to zero, their removal does not affect model convergence. In this work, we analyze whether the removal of low-magnitude non-zero gradients would help to accelerate model training with minimal performance degradation.
@@
Possible further directions and trends. Even when isolating experiments to feature scaling, there remains significant flexibility in training strategies. For example, work  (Li et al., 2022) demonstrated that training GPT-like models with short sequences initially and gradually increasing them enables stable, efficient training with larger batch sizes and learning rates. Similarly, authors of  (Di Teodoro et al., 2024) found that starting with fewer negative samples and increasing their count later boosts SRS model performance. Moreover, based on our results, an adaptive threshold to prune gradients can be employed. While these adaptive strategies are promising, we focus on studying simpler non-adaptive approaches to establish a stronger foundation for future research.
@@ 3.1. Cut-Cross-Entropy for RecSys
Our proposed method, $\mathrm{CCE}^{-}$, builds upon CCE, which itself extends the Liger-Kernel (LG) (Hsu et al., 2024a). Below, we provide a brief review of both $\mathrm{CCE}^{-}$ and LG, highlighting their key differences. We then focus on how $\mathrm{CCE}^{-}$ introduces specific improvements tailored for sequential recommendation systems.
During training, we deal with two types of memory: GPU memory (High Bandwidth Memory, HBM) and on-chip shared memory (Static Random-Access Memory, SRAM). HBM is large but relatively slow and is used to store data for computations. GPU kernel loads input data from HBM into SRAM to perform computations. SRAM is significantly faster than HBM but has much less capacity. Therefore, only small chunks of input data can be loaded into SRAM at a time. Once computations are complete, the output is saved back to HBM. While there are more details, see (Dao et al., 2022),  (Wijmans et al., 2024), this structure shapes the main goal: reduce the amount of data loaded into GPU memory. It can be achieved by minimizing the overall amount of memory used by a model or optimizing transfers between HBM and SRAM.
@@ 3.2. Sparse Gradients in Training with CE loss
In the recommender systems, the objective is to predict the next item from a predefined set $V$ of |V||V|{} items. Similar to the next-token prediction, there is exactly one correct class and |V|1|V|{}-1 incorrect classes. Thus, employing loss functions such as the CE loss provides a suitable estimate of model quality  (Di Teodoro et al., 2024).
@@ 3.3. Cut-Cross-Entropy with Negative Sampling
As demonstrated in (Klenitskiy and Vasilev, 2023), negative sampling offers dual advantages: it allows better accuracy on certain datasets and accelerates training compared to full vocabulary utilization. Motivated by this, we extend $\mathrm{CCE}^{-}$ by integrating negative sampling into it.
@@
To achieve a memory-efficient computation for this term, we developed a specialized kernel that fuses these operations. For this purpose, we adopt an approach for the online softmax computation. As a result, all intermediate operations are carried out on-chip using shared memory, while only the log-sum-exp vector is stored in the global memory (Milakov and Gimelshein, 2018; Hsu et al., 2024b).
@@
During the forward pass, the output embedding $\mathbf{E}$ is divided into the blocks $E_{n,D}$ of size $N_{B}\times D$. Since the inner dimension $D$ in transformer models for RecSys is typically not large (in  (Kang and McAuley, 2018), $D<50$; in  (Klenitskiy and Vasilev, 2023)(Klenitskiy et al., 2024), $D=64$; in  (Guo et al., 2024), $D<400$ ), division is not performed along the $D$ dimension. The matrix of the linear classifier $\mathbf{C}$ is loaded inside a for-loop by columns, determined by the block of indices $\mathbf{Inds}_{n,j}$ of size $N_{B}$.
@@ 4.1. Datasets
We use six datasets: 30Music  (Turrin et al., 2015), Zvuk  (Shevchenko et al., 2024), MovieLens-20M (ML 20M)  (Harper and Konstan, 2015), MegaMarket (Shevchenko et al., 2024), Beauty  (McAuley et al., 2015), and Gowalla  (Cheng et al., 2013). Our main focus is on datasets with a large item catalog: Zvuk, MegaMarket, Gowalla, and 30Music, which include more than 300K items. Moreover, Zvuk and MegaMarket are some of the largest open datasets in the recommendation domain  (Shevchenko et al., 2024). MovieLens and Beauty are added because of their popularity, used in more than 14 and 45 papers according to (Klenitskiy et al., 2024). For Beauty, we employ its third version, which has the largest item catalog. The characteristics of all datasets after preprocessing are in Table 2.
@@ 4.2. Evaluation Methodology
Global Temporal Cutoff: To ensure alignment with real-world sequential prediction scenarios, we use a global temporal split with the last interaction as a target and ”by user” validation scheme (Gusak et al., 2025). We split interactions at the 90% timestamp quantile. All interactions before this cutoff form the training/validation pool, while interactions after constitute the test set. Validation Set: From the training pool, we randomly select 5% of users. For each selected user, their last interaction is withheld for validation. Test Set: For users in the post-cutoff test set, only their last interaction is retained for evaluation, while all previous ones are passed as model input.
@@
Following prior work  (Anelli et al., 2019; Zhao et al., 2022; Shevchenko et al., 2024), we prioritize NDCG@10 due to its strong correlation with overall ranking quality. Along with it, record other metrics, including Coverage and Surprisal (Vasilev et al., 2024). All metrics are available in our publicly released dataset.
@@
We primarily focus on the SASRec model, as it is more commonly used in practice and frequently outperforms Bert4Rec (Klenitskiy and Vasilev, 2023; Mezentsev et al., 2024). However, to demonstrate the efficiency of CCE and $\mathrm{CCE}^{-}$ for other architectures, we also present the results of applying these loss functions to the training of BERT4Rec in Section 5.6